a guy wire is necessary to support
the telephone pole. If the height of the pole is 36 feet and
the guy wire is to be placed 15 feet from the base of the pole
how long a wire is needed?

Answer by mananth(16949)

pole height =46
horizontal distance from pole on ground to where it is pegged = 15 ft.
they form the legs of a right triangle =
let length of wire be l
..
l^2=36^2+15^2
l^2=1521
l=39 feet